UT Nov 23 2021 - D/E Night



Night Info

  • Sunset 17:20
  • Twilight Ends 18:41
  • Moonrise 19:42
  • Twilight Starts 05:31
  • Sunrise 06:52

Night Plan

As of noon, AO is now available - proceed with AO plan below first. If AO fails, and conditions are excellent - reconfigure for MODS-Bino to execute M31 Nova. If conditions are poor and AO fails, stay with LUCI for backups.

If NO AO available:

  • Configure for bino-MODS
  • SpecStd Feige 110 in twilight
  • M31 Nova observations - bin-MODS - 2.5hr - until 21:15hr
    • File Location: /home/lbto/Service_Observing/2021B_D_E/MODS/RMWagner_M31_Nova
      • m31n2008_pa4pt29_1pt2.acq
      • m31n2008_1pt2_extended1800.obs
      • Updated scripts to put bright star N of nova into target (PA = 4.29 based on SX and DX r-band acquisition images)
    • up to 0.5 mag cloud extinction, seeing up to 1.5"
    • acq image may need 180sec
    • 1800sec exposures
    • hits 30deg elevation limit at 02hr
  • Reconfigure to LBCs 21:15-21:45hr
  • LBC/TMS Testing 21:45-02:45hr
    • any conditions
  • Reconfigure to LUCIs 02:45-03:15hr
  • Dwarf Merger SDSSJ1148+2149 - 1.5hr
    • OT: /home/lbto/Service_Observing/2021B_D_E/LUCI/Rothberg_Science/Dwarf_Mergers/DwarfNergers_OT_V4.xml
    • up to ~1mag extinction, seeing up to 2"
    • Fraternal L1 zJsepc, L2 HKspec - if one LUCI is offline, then observer HKspec only
    • Telluric star first (rises before target)
    • Target hits 30deg elevation 03:33hr
    • Run script 2x (script is 30min shutter time more if time permits)
  • Backup Programs - if conditions too poor for M31 Nova
    • Start of night - MODS
      • OT File: /home/lbto/Service_Observing/2021B_D_E/PoorWeather/LocalMergers/COBandhead_Mergers_V30.xml
      • MODS LS - Arp 230 until 21:15hr - then swtich to LBCs
        • Note: There is also a LUCI LS OB for Arp 230 - that can be executed as a backup another time - OR - if already configured for LUCIs
    • If conditions are worse than limits for Dwarf Mergers by 02hr during LBC/TMS, reconfigure at 03hr to LUCIs
    • Local Mergers - NGC 4004
      • OT File: /home/lbto/Service_Observing/2021B_D_E/PoorWeather/LocalMergers/COBandhead_Mergers_V30.xml
      • LUCI LS observations of NGC 4004
      • Target hits 30deg elevation at 03:30hr, Telluric at 04hr

If AO is Available:

  • Configure for bino-LUCIs
  • AO Testing - min 0.5 night - will try to execute as much AO as possible during the night
    • thin clouds
    • 1.2" or better for full AO, 1.5" or better for ESM
    • OT files:
      • /home/lbto/Service_Observing/2021B_D_E/LUCI/AO_ESM
      • ESM is ESM/ESM_V26.xml
      • AO is AO/AO_Testing_V32.xml
        • Star on-axis testing use folder "AO STARS - FULL DITHERS *NEW*"
        • Science-like targets: Einstein's Cross is in folder "AGN/ULIRGs/LIRGs"
  • OSCO during twilight
  • Full AO testing on-axis
    • Bin 1 stars
    • Bin 2 stars
    • Bin 4 stars
  • AO science-like observations
    • ADI testing on HR8799 or IP Tau (Bin 1)
    • Einstein's Cross (Bin 4) - twilight-22:45hr
  • ESM testing
    • M52
      • Bin 1 (R=10.52) - dither sequence (off field) - 30deg limit at 01:15 - KHJ
      • Bin 4 (R=16.51 forced) - dither sequence (off field) - 30deg limit at 01:15 - KHJ
    • HCG 31
      • Bin 1 (R=10.93) - dither sequence (off field) - visible 22hr-04:30hr - KHJ
    • M51
      • Bin 4 (R=16.51 forced) - dither sequence (off field) - visible 04:30hr-twilight - Ks only
  • Test of AO w/N375 camera
  • Poor Condition Backup
    • Local Mergers:
      • OT File: /home/lbto/Service_Observing/2021B_D_E/PoorWeather/LocalMergers/COBandhead_Mergers_V30.xml
      • Switch to Arp 230 with LUCIs
        • NOTE: There is also a MODS LS OB for Arp 230 - run the one for whichever instrument is currently configured
    • Reconfigure for TMS/LBC testing at 21:15hr
    • Continue with LBC TMS testing until 02:45hr or 03hr - unless conditions improve for AO - then resume AO testing
      • Switch back to AO if 30min consistently at or better than min requirements for AO (and trends suggest no further degradation in weather beyond that)

Summary

22:40 Checking out the luci's and aligning field stops. LUCI2 mos error - strain gauges.

We are setting up with the ARGOS source to check out the AO

Juan Carlos is checking out the pupil alignment with the ARGOS Source. He has an AOS Communication issue after sending the second preset. First preset successful, but guide probe was vignetting. Second preset the AOS was stuck in acquire. Juan Carlos Rested Shell and restarted AOS.

Juan Carlos noted some master Diagnostic Communication glitch although all processes look to be up. Restarting Adsec.

Juan Carlos has noted during the daytime in the past that switchign between LBTI SOUL and LUCI SOUL there can be communication issues (new bug??)

Juan Carlos's tests successful and show that the pupil rotator is in a suitable position for observing tonight. We are set and ready! He is parking the ARGOS swing arm.

0:19 Josh is taking us down and preparing to open.

0:45 Pointed on a mag4 star. There were some left over collimation values from daytime testing.

0:47 Collimation preset. Seeing 0.65-0.75" on WFS, 0.5-1mag extinction.

0:51 Preset HD216399 for OSCO for bin1, freq 999, MODS501

0:58 OCAM Estimated mag 9.1, Rmag 7.54
TN#010059 this is NCPAoff GOPT0.42 luci1..0010-29
Offset back to the start with NCPA on and loop failed with wrong magnitude, clouds coming through just as we offset. Doug resumed through AOS when we got flux back.
RTN#010800 with NCPA. luci1...030. We opened luci1.0037

1:10 We are clouded out right now and will wait to see what we get.

1:16 Doug reclosed between the clouds. Gains optimized without Gopt3.25 0.57 0.59, with GOPT 1.23 0.22 0.22. We safeskipped and reopened.. We have 6.5 mags of extinction.

1:20 With 6.5 mags of extinction we are in a holding pattern.

1:41 We are closing up. Clouds are too thick for AO testing. We will execute close dome testing with ARGOS cal source while we wait for these to thin out.
  • Juan Carlos is going to tweak the bin4 pupils
  • NCPA check.
2:30 The bin4 pupils look highly aberrated when closed. Problem with reconstructor? Juan Carlos is not sure whats happening and was unclear about things. Perhaps see his AO log for details. He is working on the bin 4 pupils. Took several Jvan images. Moving the Guide proble moved a reflection seen in the lower right. luci1...0039-44

luci1.20211123.0039.fits 2021-11-23T02:27:24.4503 54
luci1.20211123.0040.fits 2021-11-23T02:30:18.2673 54
luci1.20211123.0041.fits 2021-11-23T02:33:35.8991 54
luci1.20211123.0042.fits 2021-11-23T02:41:04.2750 54
luci1.20211123.0043.fits 2021-11-23T02:44:38.1035 54
luci1.20211123.0044.fits 2021-11-23T02:45:24.6773 54

3:32 Clouds appear to be breaking up so Juan Carlos is parking the ARGOS source and we are going to open up. There are still some clouds but satellir looks much better.

3:45 Sending preset to HD216399 for OSCO. Rerotator in the red. Instrument rotator at -53. Need to rotate 360 and resend.

3:50 Guide star not found when preset resent. Josh is doing a pointing check but clouds are super thick again so not hopeful.

4:16. Clouds are just too thick to do anything. We are going to close up and resume closed dome AO testing.

4:32 ARGOS cal source is in place. Sending preset for 17th magnitude target to test bin4.

Changing the rerotator to optimize for correction so that the optmize gain is stressing the mirror a bit less. 1.6, 0.18, 0.36

180.25 current puptil rotator offset
moving in setips of 0.1: 180.35, 180.45, 180.55, ... no clover leaf looks ok
181.25 starting to see a little bit of clover leaf on one edge, 181.35 definitely clover leaf present

going the other direction 180.25, 180.15, 180.05, ....179.15 ... 178.65 ..178.35...
176.85 now the closer leafs on right.

So we want the midpoint to be the midpoint between those two points so 179.1 degrees.

Optimizing at 179.1, optmize gain now optimizes to 1.96, 0.18, 0.36

5:03 Bin1 full AO, perhaps a bit elongated?:

luci1.20211123.0045.fits 2021-11-23T05:03:08.4806 501
luci1.20211123.0046.fits 2021-11-23T05:06:55.4011 501
luci1.20211123.0047.fits 2021-11-23T05:22:11.4724 501

5:40. Bin1 looks flat and PSF looks ok on luci, although it does look like there might be a disturbance file on there.

5:46 Bin4 full AO:

luci1.20211123.0048.fits 2021-11-23T05:46:10.0358 54
luci1.20211123.0049.fits 2021-11-23T05:49:29.4389 54
luci1.20211123.0050.fits 2021-11-23T05:58:10.5656 54

bin4 TTM
luci1.20211123.0051.fits 2021-11-23T06:06:18.2347 13
luci1.20211123.0052.fits 2021-11-23T06:08:24.7506 13
luci1.20211123.0053.fits 2021-11-23T06:28:02.8075 13
luci1.20211123.0054.fits 2021-11-23T06:30:06.6995 13

Bin4 full AO do not look flat on OCAM and psf is very distorted. Juan Carlos tried homeing ADC's, reflattening, etc. But this persists suggesting the problem may be with reconstructor. Trying bin4 ACE
TN#055754 ADC's homed bin 4 ACE

6:00 TTM preset bin4 looks ok with modulation 20 (center pupils) but not when we ramp to modulation 3.

6:46 Full AO after homing everything looks good:

luci1.20211123.0055.fits 2021-11-23T06:46:14.2656 501
luci1.20211123.0056.fits 2021-11-23T06:49:04.2317 501
luci1.20211123.0057.fits 2021-11-23T06:54:49.9085 501
luci1.20211123.0058.fits 2021-11-23T06:56:49.9772 501

bin4 full and TTM:

luci1.20211123.0059.fits 2021-11-23T07:06:26.3926 54
luci1.20211123.0060.fits 2021-11-23T07:09:59.0010 54
luci1.20211123.0061.fits 2021-11-23T07:11:17.4642 54
luci1.20211123.0062.fits 2021-11-23T07:36:25.0679 54
luci1.20211123.0063.fits 2021-11-23T07:37:24.1251 54

luci1.20211123.0064.fits 2021-11-23T07:52:28.9569 12
luci1.20211123.0065.fits 2021-11-23T08:00:40.4975 12
luci1.20211123.0066.fits 2021-11-23T08:04:29.8288 12

8:22 bin2 full Ao looks goo:

luci1.20211123.0067.fits 2021-11-23T08:22:13.5168 250
luci1.20211123.0068.fits 2021-11-23T08:26:20.3004 250

8:24 Verified bin1 TTM and AO, bin2 full AO looks good with flat pupils. Just bin 4 is wonky.

8:34 Playing more with bin4. Juan Carlos tried the pupils the reconstructor was generated with. X & Y railed. Still doesn't look great.

9:25. All avenues are exhausted. Juan carlos is warming OCAM and goign to restart the software.

9:15 Restart of full software does not look like it helped. Pupils still looks bad and images on luci look really bad.

Bin4 full AO

luci1.20211123.0070.fits 2021-11-23T09:51:39.2678 54
luci1.20211123.0071.fits 2021-11-23T09:53:52.4709 54
luci1.20211123.0072.fits 2021-11-23T09:55:44.4345 54

Time stamps mean nothing now...

Remade the flat after cleaning things up with a bin1 corrected pupil. Saved as 20211123.sav. Pupils much flatter


Testing optimal pupil rotator positionfor abberations with bin 4. Incrementing around we found 178.25 to be mid range for where the pupils start to degrade. Doug updated the wfs conf file

Flatter pupils have helped disentabgle issues. We are still seeing abberations build as soon as we appply any ho1 to bin4.

Testing with old pupils files for bin 4 Full AO:
  • AO pupil pupil tested 20211117_192657 up until now
  • pointing back to old pupil files from 20211025
    • With the old pupil file 20200916_173857 TN#120547 Image looks round on luci, we actually have a psf
  • Pointing back to the new pupil files again for comparison at the same rotor angle of 158. The correction looks bad and opened. So that has narrowed it down, the problem is with the pupils.
-- BarryRothberg - 22 Nov 2021
Topic revision: r11 - 23 Nov 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback